हिंदी

If the letters of the word ASSASSINATION are arranged at random. Find the probability that all A’s are not coming together

Advertisements
Advertisements

प्रश्न

If the letters of the word ASSASSINATION are arranged at random. Find the probability that all A’s are not coming together

योग
Advertisements

उत्तर

Total number of word is ASSASSINATION are 13.

Where, we have 3A’s, 4S’, 2I’s, 2N’s, 1T’s and 1O’s.

If all A’s are coming together, then three are 11 alphabets

Number of words when all A’s come together

= `(11!)/(4!2!2!)`

∴ Probability when all A’s come together

= `((11!)/(4!2!2!))/((13!)/(4!3!2!2!))`

= `(11!)/(4!2!2!) xx (4!3!2!2!)/(13!)`

= `(11! xx 3!)/(13!)`

= `6/(13 xx 12)`

= `1/26`

∴ Required probability when all A’s do not come together

= `1 - 1/26 = 25/26`

shaalaa.com
  क्या इस प्रश्न या उत्तर में कोई त्रुटि है?
अध्याय 16: Probability - Exercise [पृष्ठ २९८]

APPEARS IN

एनसीईआरटी एक्झांप्लर Mathematics Exemplar [English] Class 11
अध्याय 16 Probability
Exercise | Q 14.(c) | पृष्ठ २९८

वीडियो ट्यूटोरियलVIEW ALL [1]

संबंधित प्रश्न

A coin is tossed twice, what is the probability that at least one tail occurs?


A die is thrown, find the probability of following events:

  1. A prime number will appear,
  2. A number greater than or equal to 3 will appear,
  3. A number less than or equal to one will appear,
  4. A number more than 6 will appear,
  5. A number less than 6 will appear.

Check whether the following probabilities P(A) and P(B) are consistently defined

P(A) = 0.5, P(B) = 0.7, P(A ∩ B) = 0.6


Check whether the following probabilities P(A) and P(B) are consistently defined

P(A) = 0.5, P(B) = 0.4, P(A ∪ B) = 0.8


Fill in the blank in following table:

P(A) P(B) P(A ∩ B) P(A ∪ B)
0.35 ... 0.25 0.6

A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn from the box, what is the probability that all will be blue?


If 4-digit numbers greater than 5,000 are randomly formed from the digits 0, 1, 3, 5, and 7, what is the probability of forming a number divisible by 5 when, the digits are repeated?


Two unbiased dice are thrown. Find the probability that the total of the numbers on the dice is greater than 10.

 

A bag contains 8 red, 3 white and 9 blue balls. If three balls are drawn at random, determine the probability that all the three balls are blue balls 


A bag contains 8 red, 3 white and 9 blue balls. If three balls are drawn at random, determine the probability that  all the balls are of different colours.


A bag contains 5 red, 6 white and 7 black balls. Two balls are drawn at random. What is the probability that both balls are red or both are black?


If a letter is chosen at random from the English alphabet, find the probability that the letter is a consonant .


Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(iii) 0.7 0.06 0.05 0.04 0.03 0.2 0.1

In a single throw of three dice, find the probability of getting the same number on all the three dice.


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability that all 10 are good


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is numbered 1, 2, 3, 4 or 5


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is white and numbered higher than 12 or yellow and numbered higher than 26.


Three-digit numbers are formed using the digits 0, 2, 4, 6, 8. A number is chosen at random out of these numbers. What is the probability that this number has the same digits?


Four candidates A, B, C, D have applied for the assignment to coach a school cricket team. If A is twice as likely to be selected as B, and B and C are given about the same chance of being selected, while C is twice as likely to be selected as D, what are the probabilities that A will not be selected?


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(A)


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine Probability of exactly one of the three occurs


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all the three balls are red


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all one ball is red and two balls are white


If the letters of the word ASSASSINATION are arranged at random. Find the probability that two I’s and two N’s come together


If the letters of the word ASSASSINATION are arranged at random. Find the probability that no two A’s are coming together


In a non-leap year, the probability of having 53 tuesdays or 53 wednesdays is ______.


Three numbers are chosen from 1 to 20. Find the probability that they are not consecutive ______.


The sum of probabilities of two students getting distinction in their final examinations is 1.2


C1
Probability
C2
Written Description
(a) 0.95 (i) An incorrect assignment
(b) 0.02 (ii) No chance of happening
(c) – 0.3 (iii) As much chance of happening as not
(d) 0.5 (iv) Very likely to happen
(e) 0 (v) Very little chance of happening

A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn from the box, what is the probability that atleast one will be green?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×